### v5

# Competency-Aware Planning for Probabilistically Safe Navigation Under Perception Uncertainty

Abstract
Perception-based navigation systems are useful for unmanned ground vehicle (UGV) navigation in complex terrains, where traditional depth-based navigation schemes are insufficient. However, these data-driven methods are highly dependent on their training data and can fail in surprising and dramatic ways with little warning. To ensure the safety of the vehicle and the surrounding environment, it is imperative that the navigation system is able to recognize the predictive uncertainty of the perception model and respond safely and effectively in the face of uncertainty. In an effort to enable safe navigation under perception uncertainty, we develop a probabilistic and reconstruction-based competency estimation (PaRCE) method to estimate the model’s level of familiarity with an input image as a whole and with specific regions in the image. We find that the overall competency score can accurately predict correctly classified, misclassified, and out-of-distribution (OOD) samples. We also confirm that the regional competency maps can accurately distinguish between familiar and unfamiliar regions across images. We then use this competency information to develop a planning and control scheme that enables effective navigation while maintaining a low probability of error. We find that the competency-aware scheme greatly reduces the number of collisions with unfamiliar obstacles, compared to a baseline controller with no competency awareness. Furthermore, the regional competency information is particularly valuable in enabling efficient navigation.
I INTRODUCTION
Within the area of unmanned ground vehicle (UGV) navigation in unstructured environments, learning-based components have become commonplace and seemingly necessary [1]. Deep neural network (DNN)-based perception models have proven to be particularly useful for accurately analyzing the traversability of the terrain, preventing collisions, and enabling effective navigation [1]. Unfortunately, as data-driven and black-box methods, DNN-based perception models face inherent limitations: they lack transparency and explainability [2], are often overconfident in their predictions [3], are sensitive to shifts in the input data distribution [4], and can fail quite surprisingly and ungracefully [5].
Figure 1: We develop a probabilistic reconstruction-based method for estimating the overall and regional competency of a perception model for a given input image. This competency score and competency map are used to design a sampling-based path planning algorithm that evaluates the minimum competency associated with sampled paths, removes trajectories that fall below the minimum probability threshold, and selects the best path towards the goal. An LQR-based controller is then used to track the planned path.
With the prevalence of DNN models in many real-world applications and concern about their failure modes, there has been extensive research on quantifying uncertainty in these models [6] and detecting inputs that are outside of their training distribution [7]. This has enabled the development of uncertainty-aware navigation schemes that explicitly consider the existence of uncertainty arising from DNN-based perception models. However, existing navigation schemes generally rely on uncertainty estimation methods that do not perform well for data outside of the model’s training distribution [8] or on methods that provide binary decisions about whether an input is anomalous based on the training set [9]. These methods do not allow for the design of probabilistically safe control schemes that maintain performance under multiple aspects of predictive model uncertainty.
To address these limitations, we develop a competency-aware perception, planning, and control framework to enable safe navigation in complex environments under DNN-based perception model uncertainty (Figure 1). In particular, (1) we develop an overall competency score that represents the probability that the model’s prediction is correct for a given image. (2) We propose a regional competency method that estimates the probability that regions in the environment are unfamiliar to the model. (3) We generate a method for evaluating the probability of error associated with a proposed control sequence based on the competency estimates. (4) We develop a navigation scheme that maintains efficient performance, while ensuring the probability of error remains below a user-defined error threshold. (5) We evaluate this control scheme in a challenging lunar environment with unfamiliar obstacles and demonstrate the ability of our navigation framework to successfully and efficiently navigate to goal regions, while maintaining a low probability of error.
II BACKGROUND & RELATED WORK
Field robots used in search and rescue, planetary exploration, and agricultural tasks must navigate in unstructured environments, which lack clearly viable paths or helpful landmarks [1]. In such environments, perception provides the necessary information to make the vehicle aware of the surrounding environment. Within this context, most work has focused on DNN-based methods to analyze the capability of a UGV to stably reach a terrain region and determine appropriate control actions [1]. Without properly considering the uncertainty surrounding the predictions of these DNN-based perception models, vehicles that rely on these models can face unsafe situations and system failures. This has led to a branch of uncertainty-aware navigation focused on dealing with perception model prediction uncertainty.
II-A Uncertainty Quantification (UQ)
Predictive uncertainty generally encapsulates both data/aleatoric uncertainty, which arises from complexities of the data (i.e., noise, class overlap, etc.), and model/epistemic uncertainty, which reflects the ability of the perception model to capture the true underlying model of the data [10]. The modeling of these uncertainties can generally be divided into methods based on (1) Bayesian neural networks (BNNs) that extract uncertainty as a statistical measure over the output of a BNN [11, 12], (2) deterministic neural networks (NNs), often relying on Monte Carlo (MC) dropout as approximate Bayesian inference [13], and (3) ensembles of NNs that combine the predictions of multiple deterministic networks to form a probability density function [14]. While these methods address the typical overconfidence in NN predictions with better-calibrated confidence scores, they generally focus on predictions for in-distribution inputs, which come from the same distribution as the training data. These approaches are not sufficient in general to appropriately assign confidence scores for out-of-distribution (OOD) inputs that differ significantly from those seen during training [8].
II-B Out-of-Distribution (OOD) Detection
Many recent approaches have focused on quantifying distributional uncertainty caused by a change in the input data distribution [9]. Approaches that are specifically focused on determining if an input falls outside of the input-data distribution are referred to as out-of-distribution (OOD) detection methods. These approaches are generally either (1) classification-based [15, 16, 17], (2) density-based [18, 19, 20, 21, 22], (3) distance-based [23, 24, 25, 26], or (4) reconstruction-based [27, 28, 29, 30, 31, 32, 33]. While these methods do not fully capture the predictive uncertainty associated with DNNs, they can identify inputs that are drawn from a different distribution than those used to train the model. For these inputs, the model cannot be safely trusted to obtain the correct output.
II-C Anomaly Detection & Localization
A field of study that is closely related to OOD detection is anomaly detection and localization. The task of segmenting the particular pixels containing anomalies has become popular for identifying defects in industrial inspection [34], flagging abnormalities in medical image analysis [35, 36], and detecting abnormal behavior in surveillance applications [37]. Within the area of anomaly detection and localization, most approaches are (1) reconstruction-based [38, 39, 40, 41], (2) classification-based [42, 43], or (3) distance-based [44, 45, 46, 47]. While methods in uncertainty quantification, OOD detection, and anomaly localization can successfully capture different aspects of predictive uncertainty associated with DNNs, it is not clear how these methods should be used to develop safe and robust planning and control schemes.
II-D Uncertainty-Aware Navigation
There are various approaches that seek to enable short-term, local control under perception model predictive uncertainty. Existing approaches tend to rely on one of the UQ methods discussed in Section II-A, using either BNNs [48, 49], MC dropout [50], or ensembles [51, 52, 53] to estimate uncertainty and inform the control scheme. As mentioned previously, while these methods can effectively estimate some aspects of predictive uncertainty, they are insufficient for quantifying uncertainty associated with OOD samples, which a UGV is bound to encounter when navigating previously unseen or dynamic environments. There has also been work that leverages uncertainty estimation approaches that relate more closely to the distance-based OOD detection methods [54, 55] and classification-based anomaly detection methods [56, 57]. While these approaches do not fully capture the predictive uncertainty of the perception models, they are better equipped to handle inputs that differ significantly from those seen during training.
We develop a reconstruction-based method for predictive uncertainty estimation with the end goal of probabilistically safe control of a UGV. Our approach for uncertainty estimation has several benefits compared to existing methods. (1) It captures multiple aspects of uncertainty and can effectively identify OOD or anomalous samples, unlike traditional UQ techniques. (2) In contrast to methods that simply seek to identify samples that differ significantly from the training set, our method is directly tied to the predictive uncertainty of the perception model used in a control task. Furthermore, rather than making a binary decision, as is typical in OOD and anomaly detection tasks, our method is probabilistic and can be incorporated into a probabilistically safe control scheme. (3) Finally, our method can characterize uncertainty at a regional level, rather than considering uncertainty only for the image as a whole. This offers a greater amount of model explainability and provides more flexibility in designing an uncertainty-aware planning and control scheme. We show how our method for predictive uncertainty estimation can be used to evaluate the probability of error associated with vehicle trajectories and to select a preferred trajectory based on the desired level of conservativeness. [^1] The code for reproducing our methods and results is available on GitHub: https://github.com/sarapohland/parce-nav.git.
III PROBLEM FORMULATION
In this work, we consider a UGV navigating a simulated lunar environment, in which robust visual navigation is very advantageous. During training, the perception model learns to distinguish between different regions in the environment (e.g., bumpy terrain, smooth terrain, inside crater, edge of crater, etc.) to navigate effectively. During evaluation, it encounters astronauts and human-made obstacles that were not seen during training. These unfamiliar obstacles are outlined in white in Figure 5 and are considered OOD because they were not present in the training set. The vehicle is expected to recognize regions in the environment that are unfamiliar to the perception model and continue to navigate effectively in the face of predictive uncertainty. To enable effective navigation, we develop a probabilistic competency estimation method for the perception model (Section IV), then use these competency estimates to develop competency-aware planning and control schemes (Section V). Additional details and parameters are provided in Appendices -A and -B.
IV MODEL COMPETENCY ESTIMATION
We develop a probabilistic reconstruction-based method for competency estimation and evaluate the ability of this method to distinguish between correctly classified in-distribution (ID) samples, incorrectly classified ID samples, and out-of-distribution (OOD) samples (§IV-A). We then extend this approach to estimate model competency for regions in an image and evaluate the ability to distinguish between regions in ID samples, familiar regions in OOD samples, and unfamiliar regions in OOD samples (§IV-C).
IV-A Estimating Overall Model Competency
Let $f$ be the true underlying model of the system from which our images are drawn and $\hat{f}$ be the predicted model (referred to as the perception model). For an input image, $\bm{X}$, the perception model aims to estimate the true class of the image, $f(\bm{X})$, from the set of all classes, $\mathcal{C}$. The competency of the model for this image is given by
$\rho(\bm{X}):=P\bigl{(}\{\hat{f}(\bm{X})=f(\bm{X})\}|\bm{X}\bigr{)}.$ (1)
To simplify our notation, let $\hat{c}$ be the class predicted by the perception model (i.e., $\hat{f}(\bm{X})=\hat{c}$) such that
$\rho(\bm{X})=P\bigl{(}\{f(\bm{X})=\hat{c}\}|\bm{X}\bigr{)}.$ (2)
Often, the perception model uses the softmax function to obtain an estimate of the probability $P(\{f(\bm{X})=c)\}|\bm{X})$ for each class $c\in\mathcal{C}$. However, the perception model cannot truly estimate this probability because it is limited by the data contained in the training sample. It instead estimates the probability $P(\{f(\bm{X})=c)\}|\bm{X},D)$, where $D$ is the event that the input image is in-distribution (i.e. drawn from the same distribution as the training samples). Let us then write the following lower bound on competency:
$\rho(\bm{X})\geq P\bigl{(}D\cap\{f(\bm{X})=\hat{c})\}|\bm{X}\bigr{)}.$ (3)
This lower bound can equivalently be expressed as
$\rho(\bm{X})\geq P\bigl{(}\{f(\bm{X})=\hat{c}\}|\bm{X},D\bigr{)}P(D|\bm{X}).$ (4)
We can assume that the perception model provides an estimate of the first probability. To estimate the second probability, we design an autoencoder to reconstruct the input image, training the autoencoder with the same images used to train the perception model. A holdout set is then used to estimate the distribution of the reconstruction loss for each class. It is assumed the reconstruction loss for a given class, $\mathcal{L}_{c}$, follows a Gaussian distribution with mean $\mu_{c}$ and standard deviation $\sigma_{c}$. Let $\ell(\bm{X})$ be the reconstruction loss for image $\bm{X}$. The probability this image is drawn from the same distribution as those in the training sample is given by
$P(D|\bm{X})=\sum_{c\in\mathcal{C}}P(D|\{f(\bm{X})=c\})P(\{f(\bm{X})=c\}|\bm{X}).$ (5)
Assume now that the perception model provides an accurate estimate of $P(\{f(\bm{X})=c\}|\bm{X})$. We can estimate $P(D|\{f(\bm{X})=c\})$ as the probability that the reconstruction loss corresponding to image $\bm{X}$ aligns with the bottom $N\%$ of the training images. Because $\mathcal{L}_{c}$ is a Gaussian random variable, $N$ corresponds to a z-score, $z$, and we can estimate $P(D|\{f(\bm{X})=c\})$ as $\hat{p}_{D|C}$ in the following way:
$\hat{p}_{D|C}$ $=P\bigl{(}\{\mathcal{L}_{c}>\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$ (6)
$=1-P\bigl{(}\{\mathcal{L}_{c}\leq\ell(\bm{X})-(\mu_{c}+z\sigma_{c})\}\bigr{)}$ (7)
$=1-F_{\mathcal{L}_{c}\sim\mathcal{N}(\mu_{c},\sigma_{c})}\bigl{(}\ell(\bm{X})-\mu_{c}-z\sigma_{c}\bigr{)}$ (8)
$=1-F_{Z\sim\mathcal{N}(0,1)}\left(\frac{\ell(\bm{X})-2\mu_{c}-z\sigma_{c}}{\sigma_{c}}\right)$ (9)
$=:1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right).$ (10)
Let $\hat{p}_{c}$ be the probabilistic (softmax) output of the perception model corresponding to class $c\in\mathcal{C}$. We now have the following estimate of model competency:
$\hat{\rho}(\bm{X}):=\hat{p}_{\hat{c}}\sum_{c\in\mathcal{C}}\hat{p}_{c}\left(1-\phi\left(\frac{\ell(\bm{X})-2\mu_{c}}{\sigma_{c}}-z\right)\right).$ (11)
We refer to $\hat{\rho}(\bm{X})$ as the overall competency score for image $\bm{X}$. If the model is 100% competent on the image, the probability that its prediction is correct is one. If it is entirely incompetent, the probability its prediction is correct is zero. The competency score is thus between zero and one, with higher scores corresponding to higher levels of competency [^2] Additional details about this competency score and how it is calibrated are available on arXiv: https://arxiv.org/abs/2411.16715..
IV-B Analyzing Overall Model Competency Scores
Figure 2: Distribution of overall model competency estimates for correctly classified ID samples (left), misclassified ID samples (middle), and OOD samples (right).
The distribution of competency scores for correctly classified ID images, misclassified ID images, and OOD images is shown in Figure 2. Notice that the competency score is very close to one for nearly all correctly classified samples. It is below 0.3 for the majority of OOD images and is between 0.3 and 1.0 for nearly all misclassified ID samples.
We compare our competency estimator to twelve existing methods for UQ and OOD detection. The capability of our method to distinguish between correctly classified and misclassified samples is comparable to existing methods, but our scoring method is much better at distinguishing between ID and OOD samples. Comparing our method to competing ones, the Kolmogorov–Smirnov (KS) distance between ID and OOD distributions is two times greater or more. When seeking to identify OOD samples from a set of ID samples, our method achieves an area under the receiver operating characteristic curve (AUROC) score of nearly one, and the false positive rate (FPR) is close to zero in both cases, which is an order of magnitude smaller than existing methods. In addition, we believe our scoring method is more intuitive than others, producing scores that are always between zero and one, as opposed to existing methods with no preset range. See Appendix -C for a more thorough comparison of our competency score to existing methods.
IV-C Estimating Regional Model Competency
Suppose that, in addition to estimating the overall competency score for the image as a whole, we wish to estimate a regional competency score for each segmented region in the image. Now, instead of letting $\bm{X}$ be the entire input image, $\bm{X}$ is a segmented region of the image. In this work, segmented regions are determined by the Felzenszwalb segmentation algorithm [58]. We follow roughly the same procedures to estimate the probability that each region in the input image came from the same distribution as the training samples. In this approach, rather than designing an autoecoder to reconstruct the input image, we design an image inpainting model to reconstruct a missing segment of the input image and measure the average reconstruction loss over the pixels corresponding to that image segment.
IV-D Analyzing Regional Model Competency Maps
The distribution of competency scores for regions in ID images, familiar regions in OOD images, and unfamiliar regions in OOD images is shown in Figure 3. Notice that the competency score is close to one for most of the familiar regions (in both ID and OOD images), while it is close to zero for most of the unfamiliar regions in OOD images. We can use these competency estimates to generate regional competency maps (examples in column two of Figure 4).
We compare our regional competency estimation method to seven existing methods for anomaly detection and localization. We find that our method outperforms existing methods when tasked with distinguishing between familiar and unfamiliar pixels, achieving the lowest FPRs, the highest KS distance between familiar and unfamiliar pixels in OOD images, the second highest KS distance between pixels in ID images and unfamiliar pixels in OOD images, and the second highest AUROC values. Although its performance is not significantly better than that of PaDiM [45], PaDiM produces anomaly scores in a variable range, while our method generates probabilistic scores between 0 and 1, which are more interpretable and straightforward to incorporate into a planning framework. See Appendix -D for a more thorough comparison of our regional model competency maps to existing methods.
Figure 3: Distribution of regional competency estimates for regions in ID samples (left), familiar regions in OOD samples (middle), and unfamiliar regions in OOD samples (right).
V COMPETENCY-AWARE NAVIGATION
We use the overall model competency estimates (§IV-A) and regional competency estimates (§IV-C) to design navigation schemes with varying levels of competency awareness (§V-A) and evaluate them across five scenarios (§V-B).
Figure 4: Three examples of competency-aware planning. Column 1: Input images to perception model. Column 2: Probabilistic regional competency estimates. Blue indicates high competency (around 1.0) and red indicates low competency (around 0.0). Column 3: Sampled trajectories colored by associated minimum competency score. Column 4: Trajectories deemed safe by the competency-aware planner.
V-A Competency-Aware Path Planning
Given the vehicle’s current position and goal, along with an image of the environment from the vehicle’s perspective, the path planner determines the best trajectory to take from a set of sampled paths. To obtain a diverse set of dynamically feasible position paths, the planner samples action sequences within velocity control bounds, then predicts the vehicle’s path under those actions using the estimated dynamics. The baseline planner, without any competency-awareness, selects the best path based on the progress it makes towards the goal and the vehicle orientation with respect to the goal position.
We develop five planners that utilize different competency information and strategies for responding to low competency. (1) The planner that uses only overall competency selects the best trajectory in the same way as the baseline planner when overall competency is high. If the competency score falls below some probability threshold, the planner resorts to a safe action sequence–backing up slightly then turning. (2) The turning-based planner that utilizes only regional competency information checks if regions in the vicinity of the robot are associated with low perception model competency. If those regions fall below some probability threshold, the vehicle backs up slightly and turns away from low competency regions. (3) The turning-based planner that uses both overall and regional competency checks both measures of competency before resorting to a safe response. (4) The trajectory-based planner that uses only regional competency information maps sampled paths onto the regional competency map, determines the minimum competency value associated with each path, and removes those that fall below the threshold. The planner will then select the best path among those remaining. If there is no safe path, the vehicle will back up slightly and turn away from the low competency regions. (5) The trajectory-based planner that utilizes both overall and regional competency behaves similarly but only considers regional competency when overall competency is low. Three examples of this trajectory-based planner with full competency-awareness are shown in Figure 4.
To follow the paths generated by the competency-aware path planner, we develop a reference-tracking linear-quadratic regulator (LQR) using a linearized model of vehicle dynamics. See Appendix -A for more details on our vehicle model, competency-aware path planner, and path-tracking controller.
V-B Evaluation of Navigation Performance
We evaluate the navigation capability of the full perception, planning, and control pipeline (Figure 1) with varying levels of competency awareness and different responses to low model competency. We compare (1) a baseline controller that uses no competency information, (2) a turning-based controller that uses only overall competency scores, (3) a turning-based controller that uses only regional competency maps, (4) a trajectory-based controller that uses only regional competency information, (5) a turning-based controller that uses both overall competency scores and regional competency maps, and (6) a trajectory-based controller that uses both overall and regional competency information.
Figure 5: Five scenarios used to evaluate navigation performance: (1) Drive from one side of astronaut to the other. (2) Drive from back to front of astronaut. (3) Drive up to front of habitat. (4) Navigate around both sets of ladders. (5) Navigate between two sets of ladders.
We conduct ten trials in each of the five navigation scenarios shown in Figure 5. All of these scenarios are implemented in a simulated lunar environment built on Gazebo, and all evaluations are run in real time. We evaluate navigation performance based on: (1) Success rate: percentage of trials in which the vehicle reaches its goal within the allotted time. (2) Timeout rate: percentage of trials in which the vehicle fails to reach the goal in the allotted time. (3) Collision rate: percentage of trials in which the vehicle collides at least once. (4) Navigation time: average time required to navigate to the goal (in seconds). (5) Path length: average distance traveled by the vehicle to its goal (in meters). The results of these experiments are summarized in Table I.
Competency Controller Success Timeout Collision Navigation Path
Awareness Response Rate Rate Rate Time ($s$) Length ($m$)
None N/A 62% 38% 70% 69.61 35.34
Overall Turning 10% 90% 4% 82.86 38.20
Regional Turning 90% 10% 10% 66.25 36.93
Regional Trajectory 88% 12% 8% 66.61 34.32
Both Turning 88% 12% 2% 66.42 36.62
Both Trajectory 100% 0% 4% 64.08 34.74
TABLE I: A comparison of (1) a baseline controller with no competency awareness, (2) a turning-based controller with only overall competency information, (3) a turning-based controller with only regional competency information, (4) a trajectory-based controller with only regional competency information, (5) a turning-based controller with overall and regional competency information, and (6) a trajectory-based controller with both types of competency information.
From Table I, we notice that without competency-awareness, the collision rate for the baseline controller is very high. By adding overall competency-awareness, the collision rate is significantly reduced, but the navigation time increases notably, leading to a high timeout rate. Adding regional competency information results in far higher rates of success, while maintaining low collision rates. Success rates are similar for all four controllers that use regional competency maps but are highest for the trajectory-based controller that uses both overall and regional competency information for planning. These results demonstrate the utility of perception model competency information for improving navigation performance around obstacles that are unfamiliar to the perception model. Furthermore, regional competency maps can significantly improve navigation efficiency.
Again looking at Table I, navigation times are similar across the four controllers that use regional competency maps, but the average time is slightly lower for the trajectory-based controller that uses both overall and regional information. Path lengths are also similar across the four controllers that use regional information, but they are lowest for the two trajectory-based approaches. These results demonstrate that a trajectory-based planning approach can further improve navigation efficiency in unfamiliar or changing environments.
A more nuanced evaluation of navigation schemes across different scenarios is provided in Appendix -E. To see a visual comparison of the baseline controller, the controller using only overall competency, the turning-based controller utilizing both overall and regional competency, and the trajectory-based controller utilizing both types of competency information, please view our demo video.333 A visual comparison of navigation performance across different planners is available on YouTube: https://youtu.be/mUB2MDQZObU.
VI CONCLUSIONS
We aim to improve the safety of a UGV navigating according to a perception-based system in a complex environment with obstacles that are unfamiliar to the perception model. We believe that integrating information about the predictive uncertainty of the perception model into the navigation scheme will improve the safety and performance of the system. With this idea, we propose a novel overall competency score that can distinguish between correctly classified ID samples, misclassified ID samples, and OOD samples. We expand on this approach to develop a regional competency map that can visualize regions in an image that are familiar to the perception model and those that are not. We then explore how this competency information can be integrated into a planning and control scheme for UGV navigation.
We find that, in general, competency information is useful for collision-free navigation around obstacles that are unfamiliar to the perception model. Regional competency information is particularly valuable for efficient navigation. Furthermore, we find that simply resorting to a safe action response (e.g., backing up slightly and turning) when the perception model is incompetent is often sufficient to enable collision-free navigation. However, a trajectory-based response (e.g., evaluating proposed vehicle trajectories based on regional model competency and avoiding paths through low competency regions) improves navigation efficiency. Overall, we demonstrate the utility in integrating our competency estimation method into the navigation framework.
VII LIMITATIONS & FUTURE WORK
While the integration of competency-awareness into the navigation framework greatly reduces the number of collisions with unfamiliar obstacles, collisions do still occur for a couple reasons. The field-of-view of the vehicle’s front-facing camera is quite limited, and the vehicle currently has no visual memory, so it will collide with obstacles that it cannot see. It would be useful for the vehicle to maintain some memory of competency estimates for nearby regions that are no longer visible. The competency estimators are also imperfect, and errors in competency estimates carry through the perception, planning, and control pipeline, which can result in collisions. It would be interesting to explore how these errors propagate through the navigation pipeline. As a final note, we are currently employing a relatively simple graph-based image segmentation algorithm, whose limitations can affect the performance of our regional competency estimator. Future work could explore domain-specific segmentation methods or emerging foundation models for image segmentation to enhance performance.
There are also various other avenues for future work. First, there is currently an implicit assumption that spatial features of an image, corresponding to unfamiliar obstacles, cause the reduction in perception model competency. It would be useful to develop a more general framework for understanding model competency in cases where predictive uncertainty is not due to some particular region in the image/environment. Second, the competency estimates are currently formulated for classification problems. One could generalize these formulations to be used for other tasks. In addition, this work focused on integrating competency-awareness into a local planning and control framework. It would be interesting to develop a global planner with competency-awareness. Additionally, the current response to perception model competency is to avoid regions associated with low competency. To enable continual learning within a dynamic or unfamiliar environment, it would be useful to intentionally and cautiously explore unfamiliar regions, rather than simply avoid them. It would be interesting to consider methods for safe exploration and decision frameworks to trade off between avoidance and exploration. As a final note, while robust visual navigation is critical in lunar environments, robustness to perception uncertainty is important in other domains as well. One might consider the application of this work to other areas of field robotics.